# Testing for dice control at craps

### v1

@@ 8 The likelihood ratio test
To make this more transparent, we consider an example. Suppose we want to estimate the power of the likelihood ratio test of $H_{0}:\eta=-7/495$ versus $H_{1}:\eta>-7/495$ at the alternative $\eta=0.05$. (Assume sample size $n=500$ and significance level $\alpha=0.05$.) We use $\theta=0.199803\approx\rho^{-1}(0.05)$ and generate the random sample
@@
We can carry out the analysis for both models under the original parameterizations ($\theta\in[0,1]$ in the S–S model, $\theta\in[0,0.674351]$ in the W–S model; here 0.674351 is $\rho^{-1}(13/40$). Then we reparameterize by $\eta$, the expected gain from a one-unit pass-line bet.
@@
We find that the sample-mean test outperforms the likelihood ratio test significantly, in both models, in the sense that it has higher power at various alternatives. We limited our analysis to random samples of size $n=500$, significance level $\alpha=0.05$, and the simple null hypothesis, but undoubtedly this conclusion holds more generally.
@@ 9 Conclusions
Our first conclusion is that tests based on $\overline{L}$, the sample mean of hand-length observations, are significantly more powerful than the corresponding likelihood ratio tests, so the likelihood ratio statistic can be dismissed as a potential test statistic.
@@
Our third conclusion concerns the second reparameterization and either of the two null hypotheses. Under the S–S model, the $\overline{L}$ test is less powerful than the sample proportion of pass-line wins test. But under the W–S model, the opposite is true, though to a lesser extent: the $\overline{L}$ test is more powerful than the sample proportion of pass-line wins test. And, although it is not important, the $\overline{L}$ test under the W-S model is more powerful than the $\overline{L}$ test under the S–S model. The reason this is not important is that model choice should not depend on power functions but rather on which model best fits the data to be collected. Is the shooter attempting to achieve on-axis control or control by correlation? A preliminary sample may need to be examined to determine which model is most appropriate.
@@
with help from Mathematica.
@@
$b(\theta)$ $=-34992(15+\theta),$
$c(\theta)$ $=162(2863+296\,\theta-11\,\theta^{2}),$
$d(\theta)$ $=-9(18321+1274\,\theta-629\,\theta^{2}-62\,\theta^{3}),$
$e(\theta)$ $=17775-3558\,\theta-3299\,\theta^{2}-514\,\theta^{3}-24\,\theta^{4}.$
@@
$t(3,\theta)=\frac{256+4\,\theta+\theta^{2}}{288},$
which is clearly increasing in $\theta$.
@@
Fix an integer $x\geq 3$. Suppose we can show $a_{x,0}>0$, $a_{x,1}>0$ and, for $i=2,3,\ldots,x-2$, one of the following holds:
@@
    $a_{x,i}<0$, $0<a_{x,i-1}<-a_{x,i}$, and $a_{x,i-2}>-a_{x,i}$.
  4. 4.
    $a_{x,i}<0$, $a_{x,i-1}<0$, and $a_{x,i-2}>-a_{x,i-1}-a_{x,i}$.
In case 1, the term $a_{x,i}\,\theta^{i}$ is nonnegative. In case 2, we write
@@
which is nonnegative. In case 3, we write
@@
which is nonnegative. Finally, in case 4, we write
@@
$\quad{}=(a_{x,i-2}+a_{x,i-1}+a_{x,i})\theta^{i-2}-a_{x,i-1}\,\theta^{i-1}(1-\theta)-a_{x,i}\,\theta^{i-2}(1-\theta^{2}),$
which is nonnegative. We conclude that $t(x,\theta)$ has a nonnegative derivative with respect to $\theta$ and is therefore nondecreasing.
We have verified these conditions on $x$ by computer for each $x\leq 75$, for both the S–S model and the W–S model.
@@
for all $\theta\in[0.1]$ and for positive integers $x$ sufficiently large.
@@
$\min_{\theta}c_{1}(\theta)e_{1}^{\prime}(\theta)$ $\geq 0.0394,$ $\min_{\theta}c_{1}^{\prime}(\theta)e_{1}^{\prime}(\theta)$ $\geq-0.0966,$
$\min_{\theta}c_{2}(\theta)e_{2}^{\prime}(\theta)$ $\geq-0.0000464,$ $\min_{\theta}c_{2}^{\prime}(\theta)e_{2}^{\prime}(\theta)$ $\geq 0,$
$\min_{\theta}c_{3}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.0000793,$ $\min_{\theta}c_{3}^{\prime}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.00413,$
$\min_{\theta}c_{4}(\theta)e_{4}^{\prime}(\theta)$ $\geq 0.0142,$ $\min_{\theta}c_{4}^{\prime}(\theta)e_{4}^{\prime}(\theta)$ $\geq 0.00437,$
@@
$\min_{\theta}c_{1}(\theta)e_{1}^{\prime}(\theta)$ $\geq 0.0985,$ $\min_{\theta}c_{1}^{\prime}(\theta)e_{1}^{\prime}(\theta)$ $\geq-0.168,$
$\min_{\theta}c_{2}(\theta)e_{2}^{\prime}(\theta)$ $\geq-0.000550,$ $\min_{\theta}c_{2}^{\prime}(\theta)e_{2}^{\prime}(\theta)$ $\geq 0,$
$\min_{\theta}c_{3}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.000455,$ $\min_{\theta}c_{3}^{\prime}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.00263,$
$\min_{\theta}c_{4}(\theta)e_{4}^{\prime}(\theta)$ $\geq 0.00605,$ $\min_{\theta}c_{4}^{\prime}(\theta)e_{4}^{\prime}(\theta)$ $\geq 0.00917,$

### v2

@@ 8 The likelihood ratio test
To make this more transparent, we consider an example. Under the S–S model, suppose we want to estimate the power of the likelihood ratio test of $H_{0}:\eta=-7/495$ versus $H_{1}:\eta>-7/495$ at the alternative $\eta=0.05$. (Assume sample size $n=500$ and significance level $\alpha=0.05$.) We use $\theta=0.199803\approx\rho^{-1}(0.05)$ and generate the random sample
@@
We can carry out the analysis for both models under the original parameterizations ($\theta\in[0,1]$ in the S–S model, $\theta\in[0,0.674351]$ in the W–S model; here 0.674351 is $\rho^{-1}(13/40$)). Then we reparameterize by $\eta$, the expected gain from a one-unit pass-line bet.
Table 4 shows that the sample-mean test outperforms the likelihood ratio test significantly, in both models, in the sense that it has higher power at various alternatives. We limited our analysis to random samples of size $n=500$, significance level $\alpha=0.05$, and the simple null hypothesis, but undoubtedly this conclusion holds more generally.
@@ 9 Conclusions
Our first conclusion is that tests based on $\overline{L}$, the sample mean of hand-length observations, are significantly more powerful than the corresponding likelihood ratio tests, so the likelihood ratio statistic can be dismissed as a viable test statistic.
@@
Our third conclusion concerns the second reparameterization and either of the two null hypotheses. Under the S–S model, the $\overline{L}$ test is less powerful than the sample proportion of pass-line wins test. But under the W–S model, the opposite is true, though to a lesser extent: the $\overline{L}$ test is more powerful than the sample proportion of pass-line wins test. And, although it is not important, the $\overline{L}$ test under the W–S model is more powerful than the $\overline{L}$ test under the S–S model. The reason this is not important is that model choice should not depend on power functions but rather on which model best fits the data to be collected. Is the shooter attempting to achieve on-axis control or control by correlation? A preliminary sample may need to be examined to determine which model is most appropriate, or whether a third model might better describe the data.
@@
with help from Mathematica, and (2) follows.
@@
$b(\theta)$ $=-34992(15+2\,\theta),$
$c(\theta)$ $=162(2863+786\,\theta+55\,\theta^{2}),$
$d(\theta)$ $=-9(18321+7926\,\theta+1163\,\theta^{2}+58\,\theta^{3}),$
$e(\theta)$ $=17775+11457\,\theta+2768\,\theta^{2}+298\,\theta^{3}+12\,\theta^{4}.$
@@
$t(3,\theta)=\begin{cases}\dfrac{256+4\,\theta+\theta^{2}}{288}&(\text{S--S model}),\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\dfrac{288+27\,\theta-2\,\theta^{2}}{324}&(\text{W--S model}),\end{cases}$
which is clearly increasing in $\theta\in[0,1]$.
@@
Fix an integer $x\geq 3$. Suppose we can show $a_{x,0}>0$, $a_{x,1}>0$ and, for $i=2,3,\ldots,x-2$, the following hold:
@@
    Three consecutive negative coefficients. There are no such cases.
In case 1, if $a_{x,i-1}>|a_{x,i}|$, we write
@@
which is nonnegative; if $a_{x,i-2}>|a_{x,i}|$, we write
@@
which is nonnegative. In case 2, we write
@@
$\quad{}=(a_{x,i-2}+a_{x,i-1}+a_{x,i})\theta^{i-2}-a_{x,i-1}\,\theta^{i-2}(1-\theta)-a_{x,i}\,\theta^{i-2}(1-\theta^{2}),$
which is nonnegative. We conclude that $t(x,\theta)$ has a nonnegative derivative with respect to $\theta$ and is therefore nondecreasing in $\theta$.
Under the S–S model, we have verified these conditions on $x$ by computer for each $x\leq 75$. Under the W–S model, we have done the same, but the condition $a_{x,1}>0$ fails for $x=3$, and condition 3 fails for $x=6,7,8,9$. These five cases can easily be treated separately.
@@
for all $\theta\in[0,1]$ and for positive integers $x$ sufficiently large.
@@
$\min_{\theta}c_{1}(\theta)e_{1}^{\prime}(\theta)$ $\geq 0.0394,$ $\min_{\theta}c_{1}^{\prime}(\theta)e_{1}(\theta)$ $\geq-0.0966,$
$\min_{\theta}c_{2}(\theta)e_{2}^{\prime}(\theta)$ $\geq-0.0000464,$ $\min_{\theta}c_{2}^{\prime}(\theta)e_{2}(\theta)$ $\geq 0,$
$\min_{\theta}c_{3}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.0000793,$ $\min_{\theta}c_{3}^{\prime}(\theta)e_{3}(\theta)$ $\geq-0.00413,$
$\min_{\theta}c_{4}(\theta)e_{4}^{\prime}(\theta)$ $\geq 0.0142,$ $\min_{\theta}c_{4}^{\prime}(\theta)e_{4}(\theta)$ $\geq 0.00437,$
@@
$\min_{\theta}c_{1}(\theta)e_{1}^{\prime}(\theta)$ $\geq 0.0967,$ $\min_{\theta}c_{1}^{\prime}(\theta)e_{1}(\theta)$ $\geq-0.141,$
$\min_{\theta}c_{2}(\theta)e_{2}^{\prime}(\theta)$ $\geq-0.000575,$ $\min_{\theta}c_{2}^{\prime}(\theta)e_{2}(\theta)$ $\geq 0,$
$\min_{\theta}c_{3}(\theta)e_{3}^{\prime}(\theta)$ $\geq-0.000489,$ $\min_{\theta}c_{3}^{\prime}(\theta)e_{3}(\theta)$ $\geq-0.00297,$
$\min_{\theta}c_{4}(\theta)e_{4}^{\prime}(\theta)$ $\geq-0.0104,$ $\min_{\theta}c_{4}^{\prime}(\theta)e_{4}(\theta)$ $\geq 0.0281,$
@@
$\max_{\theta}\frac{e_{2}(\theta)}{e_{1}(\theta)}\leq 0.873,\quad\max_{\theta}\frac{e_{3}(\theta)}{e_{1}(\theta)}\leq 0.843,\quad\max_{\theta}\frac{e_{4}(\theta)}{e_{1}(\theta)}\leq 0.252.$